Ferric Chloride, FeCl3
Manganese Sulfate, MnSO4
Sodium Selenite, Na2SeO3
Zinc Carbonate, ZnCO3
Cupric Sulfate, CuSO4
Potassium Iodate, KIO3

Using the mineral sources above, and the table below, formulate a mineral premix to provide the minimum NRC recommendation for iron, manganese, copper, and iodine for 90 kg pigs. Also, determine how much dilution with a carrier is required to allow the premix to be added to the diet at 0.30%.

Mineral Requirement(mg/kg)
iron 40
manganese 2
copper 3
iodine 0.14
Answer #1
Minral Requirement mg/kg Total requirement mg Mineral source used Total molar mass of the sources g/mol Molar mass of the mineral desired g/mol Mass needed of the mineral sources to satisfie the requirements * (see below for explanation)
Iron 40 40*90=3600 FeCl3 162.2 Fe= 55.85 3.6*162.2/55.85= 10.45 g of FeCl3
Manganese 2 2*90=180 MnSO4 151 Mn=54.94 1.8*151/54.94= 4.94 g MnSO4
Copper 3 3*90=270 CuSO4 159.61 Cu=63.54 2.7*159.61/63.54= 6.78 g CuSO4
Iodine 0.14 0.14*90=12.6 KIO3 214 I=126.9 0.0126*214/126.9= 0.0212 g KIO3

* Mass needed of the mineral sources to satisfie the requirements

For Fe: 3.6 g (3600 mg) of Fe are needed to satisfie the requirements for the pigs. 55.85 g of Fe are incorporated with 162.2 g of FeCl3. Then to incorporate 3.6 g of Fe Iyou will need :

55.85 g Fe -------------- 162.2 g FeCl3

3.6 g Fe--------------------- x = 3.6*162.2/55.85= 10.45 g of FeCl3

Mineral premix is composed by:

10.45 g FeCl3, 4.94 g MnSO4, 6.78g CuSO4 and 0.0212 g KIO3

Total mass is 22.19 g .

To be added as a 0.3% of a diet.

Suppose a diet is composed by 100 kg, the Premix should be 0.3 kg

Yo need to dilute the premix with 287.81 g of water (300-12.19 g)